1. The height of grade 1 pupils is are approximately normally distributed with µ = 45 inches and σ
  2. σ
  3. σ
  4. σ
  5.  = 6.
  6. If an individual pupil is selected at random, what is the probability that he or she has a height of 42 and 47?
  7. A class of 30 of these pupils is used as a sample. What is the probability that the class mean is between 42 and 47?
  8. If a pupil is selected at random, what is the probability that is taller than 46 inches?
  9. A class of 30 of these pupils is used as sample. What is the probability that the class mean is greater than 46 inches?

 



1
Expert's answer
2021-03-22T08:30:30-0400

6) "P(42<X<47)=P(\frac{42-45}{6}<Z<\frac{47-45}{6})=P(-0.5<Z<0.33)="

"=P(Z<0.33)-P(Z<-0.5)=0.3208."


7) "P(42<\bar X<47)=P(\frac{42-45}{\frac{6}{\sqrt{30}}}<Z<\frac{47-45}{\frac{6}{\sqrt{30}}})=P(-2.74<Z<1.83)="

"=P(Z<1.83)-P(Z<-2.74)=0.9633."


8) "P(X>46)=P(Z>\frac{46-45}{6})=P(Z>0.17)=1-P(Z<0.17)=0.4325."


9) "P(\bar X>46)=P(Z>\frac{46-45}{\frac{6}{\sqrt{30}}})=P(Z>0.91)=1-P(Z<0.91)=0.1814."